Mario's car travels 468 miles on 18 gallons of gas. Will 40 gallons be enough for Mario to travel 1,200 miles? Explain.
Answer by checkley77(12569)

468/18=27 miles per gallon.
40*27=1,080 miles can be driven with 40 gallons.
1,200-1,080=120 miles is to far.